OpenStudy (anonymous):

Three men went to a hotel to rent a room, the cost of the room was $30. Each man paid $10 to the bellboy and proceeded to their room. After a little while the bellboy realized that there was a special on rooms that night and the price for the men's room should have been $25. On his way to the men's room to give them back $5, he was puzzled how he was going to split $5 as he had no change. He decided he would give them each $1 and keep the remaining $2 for himself. So each man originally paid $10, but after the bellboy gave each man $1 back, each man paid $9.

OpenStudy (anonymous):

(cont.) 9 x 3 = $27 plus the $2 the bellboy put in his pocket equals $29. The original price for the room was $30. Where did the last dollar go?

OpenStudy (anonymous):

i have the answer i just wanna watch you all struggle

OpenStudy (anonymous):

if the price of the room is 25 dollars, then each man was supposed to pay 8.33 cents

OpenStudy (anonymous):

true but the $2 should be subtracted from the $27 to show what ended up being paid for the room. Or $30 - $1 - $1 - $1 - $2 = $25

ganeshie8 (ganeshie8):

services = money 25 + 2 = 30 - 1-1-1 27 = 27

OpenStudy (anonymous):

the missing dollar in the way you presented though wasn't missing. if they were to pay 8.33 a piece but were charged 9 thats 33*3 which is where the dollar is hidden!

ganeshie8 (ganeshie8):

services taken = money spent 25 + 2 = 30 - 1-1-1 27 = 27

OpenStudy (anonymous):

muahaha

ganeshie8 (ganeshie8):

i like this one lol.. seen this many times before.. but it makes me struggle everytime i try xD

OpenStudy (anonymous):

There isn't a lost dollar...the $2 should have been subtracted from 27, not added.